OpenStudy (anonymous):

5+4^2-(3*4) /2 = 7 Right?

OpenStudy (anonymous):

no it equals 15.

mathslover (mathslover):

\[\large{5+4^2 -\frac{(3*4)}{2}}\] \[\large{5+16-\frac{12}{2}}\] \[\large{5+16-6}\] \[\large{21-6=15}\]

mathslover (mathslover):

@Kae check your soln once again

OpenStudy (anonymous):

Thanks I Accidently Did 4*2 Instead!

mathslover (mathslover):

no problem also : \[\huge{\mathbb{Welcome}\space \cal{To} \space \textbf{Openstudy}}\]

OpenStudy (anonymous):

Haha! Thanks